OpenStudy (anonymous):

General Chemistry Challenge Problem: Gases Cyanogen, a highly toxic gas is 46.2% C and 53.8% N by mass. At 25 degrees Celsius and 751 torr, 1.05g of cyanogen occupies 0.500 L. A) What is the molecular formula of cyanogen. B) Predict the molecular structure C) Predict its polarity.

OpenStudy (anonymous):

Have the answers. Lets see who knows all 3.

OpenStudy (anonymous):

Medal and fan.

OpenStudy (aaronq):

a) Formula for \(\sf C_xN_y\) \(\sf PV=nRT=\dfrac{mRT}{M}\) \(\sf M_{C_xN_y}=\dfrac{mRT}{PV}=\dfrac{1.05~g*(62.36367~L~torr~K^{-1}mol^{-1})*298.15~K }{751~torr*0.5~L}=51.99~g/mol \) \(\sf M_{C_xN_y}\approx 52~g/mol=x~12~g/mol+y~14~g/mol\) Trial and error gives \(\sf C_2N_2\) The structure is probably  The structure does not have a permanent dipole moment because of symmetry - individual dipole moments caused by the lone pairs on the N terminal atoms cancel each other out.

OpenStudy (aaronq):



OpenStudy (anonymous):

Wow, nice. All 3 steps correct.